# On the Approximation Power of SiLU Networks: Exponential Rates and Depth Efficiency

### v1

@@ Abstract
This article establishes a comprehensive theoretical framework demonstrating that SiLU (Sigmoid Linear Unit) activation networks achieve exponential approximation rates for smooth functions with explicit and improved complexity control compared to classical ReLU-based constructions. We develop a novel hierarchical construction beginning with an efficient approximation of the square function $x^{2}$ more compact in depth and size than comparable ReLU realizations, such as those given by Yarotsky. This construction yields an approximation error decaying as $\mathcal{O}(\omega^{-2k})$ using networks of depth $\mathcal{O}(1)$. We then extend this approach through functional composition to establish sharp approximation bounds for deep SiLU networks in approximating Sobolev-class functions, with total depth $\mathcal{O}(1)$ and size $\mathcal{O}(\varepsilon^{-d/n})$.
Keywords: Neural network approximation, SiLU activation, Sobolev spaces, Complexity bounds.
@@ 1 Introduction
The study of approximation capabilities of neural networks is a central topic in the theory of deep learning. Since the classical results of Cybenko [cybenko1989approximation], Hornik [hornik1991approximation], and Funahashi [funahashi1989approximate], it is known that single-hidden-layer neural networks (NN) with nonlinear activation functions can approximate any continuous function on a compact domain to arbitrary accuracy. These universal approximation theorems, however, do not address efficiency, i.e., how network size relates to approximation error for specific function classes.
Barron [cheang2001penalized] introduced risk bounds for function estimation with single-hidden-layer networks, using a penalized least-squares criterion for model selection. The development of deep networks and the ReLU activation function later led to sharper approximation results. Yarotsky [yarotsky2017error] established exponential expressivity of ReLU networks for smooth functions, and Petersen and Voigtlaender [petersen2018optimal] derived optimal bounds for function classes with variable regularity. The roles of network depth and width in approximation efficiency were clarified in [telgarsky2016benefits, lu2017expressive, arora2016understanding].
In addition to ReLU, other perspectives on neural network expressivity have been developed. Daniely, in [daniely2020neural], analyzed learning mechanisms beyond over-parameterization. More recently, smooth activation functions have been reconsidered, after initially being set aside, due to their empirical performance. The SiLU (or swish) activation [elfwing2018sigmoid, ramachandran2017searching] and its variant GELU [hendrycks2016gaussian, lee2023gelu] have been adopted in modern architectures for variational surrogate model [ayena:hal-05178445].
@@
The purpose of this work is to address this gap by providing a rigorous analysis of the approximation properties of SiLU-activated networks, including convergence rates and complexity bounds, thereby situating them within the general framework of neural network approximation theory. Our work contributes to the growing literature on the expressive power of neural networks [petersen2018optimal, lu2017expressive, telgarsky2016benefits] by providing explicit construction methods with guaranteed error bounds. The results have implications for understanding why shallow or deep networks with smooth activations perform so effectively in practice and provide theoretical guidance for the design of network architectures. The constructive nature of our proofs offers potential pathways for developing specialized networks for scientific computing applications where high precision is required.
The remainder of this article is organized as follows. Section 2 provides essential background on neural network definitions and uniform approximation concepts, establishing the foundational framework for our analysis. In Section 3, we present our main contributions, beginning with the fundamental approximation of the square function using SiLU networks, which serves as the building block for subsequent results. We then develop efficient approximations for multiplication operations, arbitrary monomials, and general polynomials, demonstrating how these constructions naturally lead to recurrent neural network architectures with shared parameters. This section concludes with our main universal approximation theorems for continuous functions and for functions defined in the unit ball of Sobolev spaces, providing explicit complexity bounds. Section 4 discusses practical implications and presents experimental results validating our theoretical findings comparatively to ReLU networks.
@@ 2 Definition of neural network and uniform approximation
A neural network
A neural network of depth $m$ (i.e., $m$ layers including input and output, or (m1)(m-1) hidden layers) is a function $N:\mathbb{R}^{p}\to\mathbb{R}^{d}$ defined as the composition of affine maps and nonlinear activation functions. More precisely, the network is given by
@@
Let $f:[-B,B]^{p}\to\mathbb{R}$ be a function. We say that a function $g$ approximates $f$ with precision $\epsilon>0$ if
$\sup_{\mathbf{x}\in[-B,B]^{d}}|g(\mathbf{x})-f(\mathbf{x})|\leq\epsilon$
This inequality ensures that the approximation error is uniformly bounded by $\epsilon$ over the entire domain $B,B]^{d}$, thereby guaranteeing pointwise accuracy throughout the considered region. In the context of neural networks, this means that the network $g$ must faithfully capture the behavior of $f$ over the entire domain without exception. The complexity of such network (its depth and size) will depend on the regularity of $f$ and the required precision $\epsilon$.
@@
Figure 1: Approximation of the function $\log(7+x)\cos(x^{3})$ by ReLU neural networks (left, Mse =1.69) and SiLU neural networks (right, MSE=1.63).
2.1 Motivations
The choice of activation function is a critical factor in the ability of neural networks to approximate complex functional behaviors. A comparison between two identical architectures ($12$ hidden layers and one linear output layer with 64 neurons per hidden layer) using the SiLU activation function and the ReLU activation function ($\max(0,x)$), respectively, reveals experimentally that the SiLU version provides superior approximation accuracy for equal complexity.
To illustrate this, consider the function
$T_{1}(x)=\log(7+x)\cdot\cos(x^{3}),$
defined on a compact interval. The network using SiLU produces a lower approximation error and yields a smoother curve, more closely matching the local variations of $f$. This superiority is explained by several fundamental properties of the SiLU function: its regularity ($C^{\infty}$), its differentiability over its entire domain, and a better-behaved gradient specifically, non-zero over $\mathbb{R}$, unlike ReLU.
These observations motivate the design of compact and optimized neural networks that specifically leverage SiLU activation for function approximation. Such an approach could be to:
  • Reduce the size of networks required to achieve a given precision,
  • Improve generalization capabilities and numerical stability,
  • Achieve enhanced performance on highly regular functions.
3 Our theoritical results
3.1 SiLU is robust for polynomial approximation
The following theorem establishes the ability of SiLU neural networks to efficiently approximate the square function, which serves as a fundamental building block for approximating arbitrary polynomials. This result demonstrates the robustness of SiLU activation for polynomial approximation, as it achieves exponential convergence with constant-depth networks.
@@ 3.1 SiLU is robust for polynomial approximation
For any $B>0$ and target accuracy $\epsilon\in(0,1)$ , there exist constants $C_{1}>0$, $\omega_{1}>1$ and a family of feedforward SiLU neural networks $\{Q_{k}\}_{k\in\mathbb{N}}$, with one hidden layer with $3$ neurons each, such that for $k\geq\max\left\{\left\lceil\frac{\ln C_{1}-\ln\epsilon}{2\ln\omega_{1}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{1}}\right\rceil\right\}$
$\sup_{|x|\leq B}|Q_{k}(x)-x^{2}|\leq\epsilon,$
@@
The complete proof is provided in Appendix A.1. We then construct the function
$g_{a}(x)=(x+a)\sigma(x+a)+(-x+a)\sigma(-x+a)-2a\sigma(a),$
which eliminates the odd-order terms, leaving only even-order terms starting from $x^{2}$. The dominant coefficient of $x^{2}$ is $K(a)=2\sigma^{\prime}(a)+a\sigma^{\prime\prime}(a)$, which is non-zero for an appropriate choice of $a$.
We normalize $g_{a}(x)$ to isolate the $x^{2}$ term and express the residual error as a series of higher even powers. To reduce this error, we introduce a sequence of functions $Q_{k}$ defined recursively via a scaling transformation:
$Q_{k}(x)=\beta^{-2k}f_{a}(\beta^{k}x),$
with $\beta<1$. This construction exponentially compresses the error terms, as the coefficients of higher-order terms are multiplied by $\beta^{k(j-2)}$ for the term in $x^{j}$.
Over any interval [B,B][-B,B], the error $|Q_{k}(x)-x^{2}|$ is bounded by a sum of terms that decay exponentially with $k$, ensuring uniform convergence with an error of $\mathcal{O}(\beta^{2k})$. Finally, we show that the function $Q_{k}$ can be implemented by a SiLU neural network of constant depth and size, using a linear combination of three SiLU evaluations with carefully chosen parameters
$Q_{k}(x)=\gamma\cdot SiLU(Wx+b)$ (1)
@@
Figure 2: Architecture of NN $Q_{k}$ (Theorem 1).
Figure 3: Approximation of the function $x^{2}$ by the functions $Q_{k}$ (Theorem 1).
@@
$xy=\frac{1}{4}[(x+y)^{2}-(x-y)^{2}]$
to establish the ability of neural networks with SiLU activation functions to efficiently approximate the multiplication of two variables. It demonstrates that the approximation error decreases exponentially fast as the complexity of the network increases, all while using an architecture whose size and depth remain constant, independent of the required level of precision.
@@
For every $B>0$ and target accuracy $\epsilon\in(0,1)$, there exist constants $C_{2}>0$, $\omega_{2}>1$ and a family of feedforward SiLU neural networks $\{M_{k}\}_{k\in\mathbb{N}}$, with at most one hidden layer with $4$ neurons each, such that for $k\geq\max\left\{\left\lceil\frac{\ln C_{2}-\ln\epsilon}{2\ln\omega_{2}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{2}}\right\rceil\right\}$
$\sup_{|x|,|y|\leq B}|M_{k}(x,y)-xy|\leq\epsilon$
@@
The complete and detailed proof can be found in Appendix A.2. The proof is built upon the classical algebraic identity that decomposes the product into a difference of squares: $xy=\frac{1}{4}[(x+y)^{2}-(x-y)^{2}]$. The neural network $M_{k}$ is explicitly constructed to implement this identity using approximate squaring units.
The construction proceeds in three conceptual steps. First, the input pair $(x,y)$ is transformed linearly into the sum $(x+y)$ and the difference (xy)(x-y). Each of these two terms is then processed by a separate SiLU network, $Q_{k}$, which approximates the square function with an exponentially decaying error of $\mathcal{O}(\omega^{-2k})$, as per Theorem 1. Finally, the outputs of these two subnetworks are combined linearly to produce the final approximation of the product. This construction leads to an explicit formula for the entire network $M_{k}$, which can be expressed as a single, shallow SiLU network with a specific weight structure
$M_{k}(x,y)=\gamma\cdot\text{SiLU}(Wz+b)$ (2)
where $z=[x,y]^{\top}$ is the input vector. The parameters are meticulously chosen to encode the entire approximation process. The weight matrix
$W=\beta^{k}\begin{bmatrix}+1&+1\\ -1&-1\\ +1&-1\\ -1&+1\end{bmatrix}$
computes the four required linear combinations $(x+y)$, (x+y)-(x+y), (xy)(x-y), and (xy)-(x-y). The bias vector $b=[a,a,a,a]^{\top}$ is a constant. The output weight vector
$\gamma=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}+1&+1&-1&-1\end{bmatrix}^{\top}$
performs the final linear combination, effectively calculating $\frac{1}{4}[Q_{k}(x+y)-Q_{k}(x-y)]$.
The exponential error rate for the product, $C_{2}\cdot\omega_{2}^{-2k}$, is a direct consequence of the error bounds from the two $Q_{k}$ subnetworks. Crucially, the entire computation is encapsulated in a network with constant depth, $\mathcal{O}(1)$, and constant size, $\mathcal{O}(1)$, as the two squaring networks are computed in parallel within this single layer. ∎
@@
Figure 5: Approximation $M_{k}\left(\cos(x),\sin(x)\log(x^{2})\right)$ of the product of the functions $\cos(x)$ and $\sin(x)\log(x^{2})$ by the NN $\{M_{k}\}$ (Theorem 2).

### v2

# Approximation with SiLU Networks: Constant Depth and Exponential Rates for Basic Operations

@@ Abstract
We present SiLU network constructions whose approximation efficiency depends critically on proper hyperparameter tuning. For the square function $x^{2}$, with optimally chosen shift $a$ and scale $\beta$, we achieve approximation error $\varepsilon$ using a two-layer network of constant width, where weights scale as $\beta^{\pm k}$ with $k=\mathcal{O}(\ln(1/\varepsilon))$. We then extend this approach through functional composition to Sobolev spaces, we obtain networks with depth $\mathcal{O}(1)$ and $\mathcal{O}(\varepsilon^{-d/n})$ parameters under optimal hyperparameters settings. Our work highlights the trade-off between architectural depth and activation parameter optimization in neural network approximation theory.
Keywords:
Neural network approximation , SiLU activation, Sobolev spaces, Complexity bounds
@@ 1 Introduction
The study of approximation capabilities of neural networks is a central topic in the theory of deep learning. Since the classical results of Cybenko [4], Hornik [9], and Funahashi [7], it is known that single-hidden-layer neural networks (NN) with nonlinear activation functions can approximate any continuous function on a compact domain to arbitrary accuracy. These universal approximation theorems, however, do not address efficiency, i.e., how network size relates to approximation error for specific function classes.
Barron [3] introduced risk bounds for function estimation with single-hidden-layer networks, using a penalized least-squares criterion for model selection. The development of deep networks and the ReLU activation function later led to sharper approximation results. Yarotsky [16] established exponential expressivity of ReLU networks for smooth functions, and Petersen and Voigtlaender [13] derived optimal bounds for function classes with variable regularity. The roles of network depth and width in approximation efficiency were clarified in [15, 12, 1].
In addition to ReLU, other perspectives on neural network expressivity have been developed. Daniely, in [5], analyzed learning mechanisms beyond over-parameterization. More recently, smooth activation functions have been reconsidered, after initially being set aside, due to their empirical performance. The SiLU (or swish) activation [6, 14] and its variant GELU [8, 11] have been adopted in modern architectures for variational surrogate model [2].
@@
The purpose of this work is to address this gap by providing a rigorous analysis of the approximation properties of SiLU-activated networks, including convergence rates and complexity bounds, thereby situating them within the general framework of neural network approximation theory.
The remainder of this article is organized as follows. Section 2 provides essential background on neural network definitions and uniform approximation concepts, establishing the foundational framework for our analysis. In Section 3, we present our main contributions, beginning with the fundamental approximation of the square function using SiLU networks, which serves as the building block for subsequent results. This section contains several key contributions:
  • With weights scale as $\mathcal{O}(\ln(1/\varepsilon))$, Theorem 1 demonstrates that the square function $x^{2}$ can be approximated with exponential accuracy $\mathcal{O}(\omega^{-2k})$ by a shallow SiLU network with only 3 neurons and constant depth. Theorem 2 extends this to multiplication $x\cdot y$ , requiring just 4 neurons in a single hidden layer.
  • Theorem 3 generalizes to arbitrary monomials $x^{m}$, showing they can be approximated with networks of depth $\mathcal{O}(m)$ and constant width.
  • Along similar lines, Theorem 4 presents an alternative single-hidden-layer construction using finite differences, achieving monomial approximation with $m+1$ neurons.
  • Theorem 5 combines these results to approximate arbitrary polynomial functions.
  • Corollary 1 reveals that the monomial approximation networks naturally form a recurrent neural network (RNN) with shared parameters, efficiently generating polynomial bases $\{1,x,x^{2},\ldots,x^{m}\}$ through sequential computation.
  • For Continuous function approximation, Theorem 6, Proposition 1 and 2 and Corollary 2 develop a step-function approach, with constant depth and size $N=\mathcal{O}\left(\frac{b-a}{\omega_{f}^{-1}(\varepsilon)}\right)$ depending on the modulus of continuity $\omega_{f}$.
  • Theorem 8 establishes optimal approximation rates for functions in Sobolev spaces $\mathcal{F}_{n,d}$, showing that SiLU networks achieve error $\varepsilon$ with size $\mathcal{O}(\varepsilon^{-d/n})$ and constant depth, eliminating the logarithmic factor present in comparable ReLU constructions.
Section 4 discusses practical implications and presents experimental validation of our theoretical results. We analyze the optimization of key parameters ($\beta$, $k$, and $a$) in our constructions and provide comparative experiments between SiLU and ReLU networks.
Our work contributes to the growing literature on the expressive power of neural networks [13, 12, 15] by providing explicit construction methods with guaranteed error bounds. The results have implications for understanding why deep networks with smooth activations perform so effectively in practice and provide theoretical guidance for the design of network architectures. The constructive nature of our proofs offers potential pathways for developing specialized networks for scientific computing applications where high precision is required.
@@ 2 Definition of neural network and uniform approximation
Definition
A neural network of depth $m$ (i.e., $m$ layers including input and output, or (m2)(m-2) hidden layers) is a function $N:\mathbb{R}^{p}\to\mathbb{R}^{d}$ defined as the composition of affine maps and nonlinear activation functions. More precisely, the network is given by
@@
Let $f:[-B,B]^{p}\to\mathbb{R}$ be a function. We say that a function $g$ approximates $f$ with precision $\varepsilon>0$ if
$\sup_{\mathbf{x}\in[-B,B]^{d}}|g(\mathbf{x})-f(\mathbf{x})|\leq\varepsilon.$
@@
The notation $h(\varepsilon)=\underset{\varepsilon\to 0}{\mathcal{O}}(1)$ means that $h$ does not depend on the $\varepsilon$, it remains bounded as $\varepsilon\to 0$. In the following, we simply write $\mathcal{O}(1)$ with $\varepsilon\to 0$ being clear from context.
Figure 1: Approximation of the function $\log(7+x)\cos(x^{3})$ by ReLU neural networks (left, Mse =1.75) and SiLU neural networks (right, MSE=1.69).
While deep ReLU networks achieve optimal worst-case approximation rates for Sobolev spaces [16], empirical evidence suggests that smooth activations like SiLU can sometimes achieve lower error on specific regular functions. For instance, in approximating $\log(7+x)\cdot\cos(x^{3})$ with identical architectures, we observe significantly lower error using SiLU compared to ReLU (Fig. 1). This observation motivates our investigation into whether smooth activations might offer improved constants or more favorable depth–width trade-offs for functions with particular smoothness or algebraic structure, complementing ReLU’s asymptotic optimality for general function classes.
To systematically investigate these potential advantages, we develop a theoretical framework for analyzing SiLU network approximation. Our analysis operates within an extended model where activation parameters can be precision-dependent—acknowledging that practical implementations often tune such parameters. Within this framework, we derive explicit complexity bounds and compare them with established ReLU constructions. This allows us to precisely characterize trade-offs between architectural depth, parameter count, and numerical precision requirements.
3 Theoretical results: Efficient polynomial approximation with SiLU
3.1 A constant-depth building block: Approximating the square function
In this section, we establish the theoretical foundations for this efficiency. We begin by showing how SiLU networks approximate the most basic nonlinear monomial, the square function $x^{2}$, with constant depth and exponential accuracy, where weights scale as $\beta^{\pm k}$ with $k=\mathcal{O}(\ln(1/\varepsilon))$ (Fig 2).
Our construction will serve as the fundamental building block for all subsequent approximations.
@@ 3.1 A constant-depth building block: Approximating the square function
For any $B>0$ and target accuracy $\varepsilon\in(0,1)$ , there exist constants $C_{1}>0$, $\omega_{1}>1$ and a family of feedforward SiLU neural networks $\{Q_{k}\}_{k\in\mathbb{N}}$, with one hidden layer with $3$ neurons each, such that for $k\geq\left\lceil\frac{\ln C_{1}-\ln\varepsilon}{2\ln\omega_{1}}\right\rceil$
$\sup_{|x|\leq B}|Q_{k}(x)-x^{2}|\leq\varepsilon,$
@@
Let $\sigma(x)=(1+e^{-x})^{-1}$ and $\text{SiLU}(x)=x\sigma(x)$. Define $g_{a}(x)=\text{SiLU}(x+a)+\text{SiLU}(-x+a)-2\text{SiLU}(a)$.
Expanding $\sigma(z)$ around $z=a$:
$\sigma(z)$ $=\sigma(a)+\sigma^{\prime}(a)(z-a)+\frac{\sigma^{\prime\prime}(a)}{2}(z-a)^{2}+\cdots$
Then, using $s=\sigma(a)$, $s^{\prime}=\sigma^{\prime}(a)$, $s^{\prime\prime}=\sigma^{\prime\prime}(a)$, and
$(a\pm x)\sigma(a\pm x)$ $=as\pm(s+as^{\prime})x+\Big(s^{\prime}+\frac{as^{\prime\prime}}{2}\Big)x^{2}\pm\cdots+,$
We have
$g_{a}(x)=(2s^{\prime}+as^{\prime\prime})x^{2}+\sum_{j=2}^{\infty}c_{2j}(a)x^{2j},$
where $c_{2j}(a)=\frac{2\bigl(s^{(2j-1)}/(2j-1)!+as^{(2j)}/(2j)!\bigr)}{2s^{\prime}+as^{\prime\prime}}$.
Assuming $K(a)=2s^{\prime}+as^{\prime\prime}\neq 0$, set
$f_{a}(x)=g_{a}(x)/K(a)=x^{2}+\sum_{j=2}^{\infty}c_{2j}(a)x^{2j}.$
Define iteratively for $k\geq 0$ with $\beta\in(0,1)$:
$Q_{0}(x)=f_{a}(x),\qquad Q_{k}(x)=\beta^{-2}Q_{k-1}(\beta x)=\beta^{-2k}f_{a}(\beta^{k}x).$
Hence
$Q_{k}(x)=x^{2}+\sum_{j=2}^{\infty}c_{2j}(a)\beta^{k(2j-2)}x^{2j}.$
A gives
$\Bigl|\frac{\sigma^{(m)}(a)}{m!}\Bigr|\leq C_{a}\,\rho^{\,m},\qquad\rho=\frac{2}{\pi},\;C_{a}>0.$
Consequently,
$|c_{2j}(a)|\leq\frac{2C_{a}(1+|a|\rho)}{|K(a)|}\,\rho^{\,2j-1}=:\widetilde{C}_{a}\,\rho^{\,2j}.$
For $|x|\leq B$,
$|Q_{k}(x)-x^{2}|$ $\leq\sum_{j\geq 2}|c_{2j}(a)|\,\beta^{k(2j-2)}B^{2j}$
$\leq\widetilde{C}_{a}\rho^{4}B^{4}\beta^{2k}\sum_{j\geq 2}(\rho^{2}B^{2}\beta^{2k})^{(j-2)}$
$=\widetilde{C}_{a}\,\rho^{4}B^{4}\,\frac{\beta^{2k}}{1-\rho^{2}B^{2}\beta^{2k}}\quad\text{if }\rho^{2}B^{2}\beta^{2k}<1.$
Choosing $\beta$ small enough that $\rho^{2}B^{2}\beta^{2}\leq\frac{1}{2}$, we obtain for all $k\geq 1$
$|Q_{k}(x)-x^{2}|\leq 2\widetilde{C}_{a}\,\rho^{4}B^{4}\beta^{2k}=C_{1}\beta^{2k},$
with $\beta^{-1}=\omega_{1}$.
Choosing $k\geq\bigl\lceil\frac{\ln C_{1}-\ln\varepsilon}{2\ln\omega_{1}}\bigr\rceil=\mathcal{O}\Bigl(\ln 1/\varepsilon\Bigr)$ ensures $|Q_{k}(x)-x^{2}|\leq\varepsilon$ for all $|x|\leq B$. Indeed,
Figure 2: Architecture of NN $Q_{k}$ (Theorem 1).
$Q_{k}(x)=\gamma\cdot\text{SiLU}(Wx+b)$ (1)
@@
The theoretical convergence shown above is confirmed numerically in (Fig. 3), which displays the approximation of the function $x^{2}$ by the NN $Q_{k}$ with $a=0$ and $\beta=0.27$(Fig. 2).
Figure 3: Approximation of the function $x^{2}$ by the NN $Q_{k}$ with $a=0$ and $\beta=0.27$.
3.2 Efficient Multiplication via Squaring
@@
$xy=\frac{1}{4}[(x+y)^{2}-(x-y)^{2}]$ (2)
to establish the ability of neural networks with SiLU activation functions to efficiently approximate the multiplication of two variables.
@@
For every $B>0$ and target accuracy $\varepsilon\in(0,1)$, there exist constants $C_{2}>0$, $\omega_{2}>1$ and a family of feedforward SiLU neural networks $\{M_{k}\}_{k\in\mathbb{N}}$, with at most one hidden layer with $4$ neurons each, such that for $k\geq\left\lceil\frac{\ln C_{2}-\ln\varepsilon}{2\ln\omega_{2}}\right\rceil$
$\sup_{|x|,|y|\leq B}|M_{k}(x,y)-xy|\leq\varepsilon$
@@
Using the identity $xy=\frac{1}{4}[(x+y)^{2}-(x-y)^{2}]$, we construct an approximant $M_{k}(x,y)$ from the square-approximating network $Q_{k}$ of Theorem 1. For $|x|,|y|\leq B$, we have $|x\pm y|\leq 2B$, so Theorem 1 guarantees
$\sup_{|x|,|y|\leq B}|Q_{k}(x\pm y)-(x\pm y)^{2}|\leq C_{1}\omega_{1}^{-2k}$
with constants $C_{1}>0,\ \omega_{1}>1$. Hence
$\sup_{|x|,|y|\leq B}|M_{k}(x,y)-xy|$ $\leq\frac{1}{4}\bigl(|Q_{k}(x+y)-(x+y)^{2}|+|Q_{k}(x-y)-(x-y)^{2}|\bigr)$
$\leq C_{2}\omega_{1}^{-2k},$
where $C_{2}=C_{1}/2$ and $\omega_{2}=\omega_{1}$.
Choosing $k\geq\left\lceil\frac{\ln C_{2}-\ln\varepsilon}{2\ln\omega_{2}}\right\rceil$ yields an error $\leq\varepsilon$. Since $Q_{k}$ has constant depth and size, and the linear operation adds only $\mathcal{O}(1)$ layer, the total network $M_{k}$ also has constant depth and size.
Explicitly, $M_{k}$ can be written as a SiLU network with weight matrix
$W=\beta^{k}\begin{bmatrix}+1&+1\\ -1&-1\\ +1&-1\\ -1&+1\end{bmatrix},\quad\gamma=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}1&1&-1&-1\end{bmatrix}^{\top}$,
bias $b=(a,a,a,a)^{\top}$, and input $z=(x,y)^{\top}$. ∎
Experimentally, we employ the neural network $\{M_{k}\}$ (whose architecture is shown in (Fig. 4) with parameters $a=0$ and $\beta=0.27$ to approximate the product $\cos(x)\cdot\sin(x)\log(x^{2})$ via $M_{k}\bigl(\cos(x),\sin(x)\log(x^{2})\bigr)$. The results are presented in Fig. 5.
@@
Figure 5: Approximation $M_{k}\left(\cos(x),\sin(x)\log(x^{2})\right)$ of the product of the functions $\cos(x)$ and $\sin(x)\log(x^{2})$ by the NN $\{M_{k}\}$ with $a=0$ and $\beta=0.27$.
3.3 Approximation of Monomials
3.3.1 Deep Network Construction with Constant Width